OpenStudy (anonymous):

a aluminum wire of unknown length has a diameter of 0.25 inch and resistance of 0.28 ohms. by several successive passes through drawing dies the diameter of the wire is reduce by 0.05inch. calculate the resistance of the reduced size wire.

OpenStudy (anonymous):

Use formula: Resistance R = resistivity* length /area Here, resistivity and length are not changed, so R is inversely proportional to area i.e radius ^2 Hope this helps.

OpenStudy (anonymous):

\[R=\frac{ PL }{A ^2 } ? do i need \to compute for the CM=(D)^2\]

OpenStudy (anonymous):

R = pl/pi*r^2, right ?

OpenStudy (anonymous):

it doesnt need to convert inch to mil?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

ok ...

OpenStudy (anonymous):

there are two diameter ... 0.25-0.05? (inch) do it need to subtract?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Convert the diameters to radii. Tell me the radii.

OpenStudy (anonymous):

No, radius is half of diameter.

OpenStudy (anonymous):

is it 0.125

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

so whats next?

OpenStudy (anonymous):

how about the lenght?

OpenStudy (anonymous):

shailkumar .. whats next?

OpenStudy (anonymous):

Ok. As you know that p and l are not change, so pl is also not change i.e pl = constant call it k

OpenStudy (anonymous):

So we have R = k/pi*r^2

OpenStudy (anonymous):

Put r = 0.125 and R = 0.28 ohm and find k

OpenStudy (anonymous):

Once you've got k, use R = k/pi*r^2 again, put r = 0.05 and the value of k you got.This will give you R

OpenStudy (anonymous):

K= 0.0137?

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

R= 1.74433 ohms?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

that's the final answer?

OpenStudy (anonymous):

What do you think? What else is left?

OpenStudy (anonymous):

none.. THANK YOU SO MUCH